# RaffeSDG: Random Frequency Filtering enabled Single-source Domain Generalization for Medical Image Segmentation

### v2

@@ III-B2 Homologous sample blending
III-C Generalizable segmentor using structural saliency
The data augmentation promises RaffeSDG a substantial increase in domain diversity, even within a single-source domain, facilitating the generalizability to unseen domains. Motivated by the inherent ability of the Gaussian filter to extract structural saliency, we incorporate robust representations into the segmentor of RaffeSDG to boost the robustness of representations.
The residual obtained by subtracting the Gaussian filter result from an image serves as a valuable indicator of its structural saliency [22]. Based on this insight, we incorporate a Gaussian filter within the segmentor to boost the learning of robust representations by structural saliency. For an image $x$, the structure saliency map $\psi(x)$ is acquired by:
@@ III-C Generalizable segmentor using structural saliency
where $g$ is a Gaussian filter with the kernel of $(r,\sigma)$. $r$ and $\sigma$ are determined according to the image size following [22].
Subsequently, the structure saliency map serves as a self-supervised pretext task for the segmentor. The self-supervision loss is defined as:
@@
Reconstructing the structure saliency map from augmented samples may promote domain-invariant representation learning. Therefore, to optimize the model through both self-supervision and segmentation supervision, we propose a coupling network as the segmentor, as illustrated in Fig. 2 (c). Specifically, the segmentor consists of an encoder and two decoders. The encoder $\mathcal{E}$ processes the augmented samples $\bar{x}_{k}$ and shares its outputs with both the decoder $\mathcal{D}_{sel}$ and $\mathcal{D}_{seg}$, where $\mathcal{D}_{sel}$ predicts the structure saliency and $\mathcal{D}_{seg}$ predicts the segmentation mask.
The concrete implementation of the segmontor is outlined in Algorithm 1. $\mathcal{E}$ is skip-connected to $\mathcal{D}_{sel}$ to facilitate the flow of contextual information, and attention mechanisms are employed to properly integrate the self-supervision into the segmentation task. The $l^{th}$h layer of the encoder and decoders are denoted as $\mathcal{E}^{l}$, $\mathcal{D}^{l}_{sel}$, and $\mathcal{D}^{l}_{seg}$, with corresponding outcomes $f^{l}_{\mathcal{E}}$, $f^{l}_{sel}$, and $f^{l}_{seg}$. The terms $CA$ and $SA$ represent typical channel and spatial attention respectively.
@@
where $\alpha$ balances $\mathcal{L}_{consist}$ and $\mathcal{L}_{seg}$.
IV Experiments
@@
IV-A Experiment settings
TABLE I: Experimental data summary
@@ IV-A Experiment settings
Vessel Fundus DRIVE, AVRDB, 40, 100, 28, 20
photography CHASEDB, STARE
@@
Joint Ultrasound 4 clinical datasets 1597, 956, 982,
@@
IV-A1 Datasets
The experiments utilized four medical imaging modalities obtained from three different human tissues, as summarized in Table I. The data for the vessel and retinal layer were publicly available, while the data for the joint cartilage were collected in collaboration with our partner, Southern University of Science and Technology Hospital, using discrepant ultrasound configurations of gain and grayscale. In order to generalize and assess segmentation models across these datasets, the ground-truth annotations were standardized. The datasets of DRIVE, HC-MS, and A were respectively employed as the single-source domain for each tissue, while the remaining datasets were treated as unseen target domains.
@@ IV-A2 Implementation
IV-B Data augmentation configuration in RaffeSDG
@@
IV-B1 Filter types and blending strategies
To validate the configuration of the proposed data augmentation approach in RaffeSDG, we explored different filter types and blending strategies. The filtered and blended samples are illustrated in Fig. 4. Moreover, training data is augmented from DRIVE, and a vanilla U-Net is trained to quantitatively evaluate the performance of various configurations on the target domains, as summarized in Table II.
@@ IV-B1 Filter types and blending strategies
Filter type: To analyze the impact of filter types on data augmentation, we employ both the ideal filter and the Butterworth filter to introduce variations into the LFS of images. Fig.4 (b) presents the function diagram and the filtered result of the ideal filter. The sharp cutoff frequency of the ideal filter leads to the presence of ringing artifacts in the filtered result. On the other hand, the Butterworth filter offers a smoother transition between the passband and stopband, reducing the occurrence of ringing artifacts. Thus, the Butterworth filter is deemed more suitable for data augmentation, as illustrated in Fig.4 (c). Additionally, to enhance diversity in the augmented samples, we implement a channel-wise filter by applying Butterworth filters with different parameters to the RGB channels, as depicted in Fig. 4 (d). Consequently, superior performance is achieved by the channel-wise filter in Table II.
Blending strategy: To interpret the effect of blending, we compare three strategies, including patch, grid, and continuous mask, as illustrated in Fig.4 (e). The images blended using patch and grid masks exhibit a noticeable gap, which deviates from real-world scenarios and hampers the overall context. In contrast, the continuous mask enables a seamless context in the augmented image, resulting in a more coherent representation that aligns well with real-world scenarios. The evaluation results provided in TableII further validate our choice of filter and blending configurations. Accordingly, the channel-wise filter and the continuous mask are cooperated to implement RaffeSDG.
TABLE II: Augmentation configuration comparison on training U-Net using only DRIVE for out-of-domain segmentation.
@@ IV-B2 Parameter range
IV-C Augmentation comparison
The comparison includes both commonly used augmentation techniques and augmentation approaches designed in SDG algorithms. Specifically, the augmentation techniques include geometric transformation, CutOut [20], and CutMix [21]. And the SDG augmentation approaches are collected from ACVC [12], SLAug [13], GIN-IPA [14], and FreeSDG [15]. In the comparison, the dataset of DRIVE is utilized as the single-source domain, and a vanilla U-Net serves as the unified segmentor, which is trained with augmented data and conducts out-of-domain inference. The augmented samples are visualized in Fig. 7, while the segmentation performance is quantified in Table III by training a U-Net.
Figure 7: Samples augmented by various augmentation techniques.
As depicted in Fig. 7, the native augmentation strategy of geometric transformation, CutOut [20], and CutMix [21] are constrained to the single-source domain, leading to limited appearance variations in the augmented samples. In contrast, the augmentation approaches used in SDG tasks present remarkable performance in Table III. For implementing SDG in classification, ACVC [12] introduces 19 visual corruptions and 3 Fourier-based corruptions to augment the single-source domain. For medical image segmentation, SLAug [13] introduces a saliency-balancing fusion strategy that combines global-scale and local location-scale aspects to augment the single-source domain. Moreover, GIN-IPA [14] proposes a causality-inspired data augmentation approach, which utilizes global intensity non-linear augmentation to diversify the appearances of medical images and addresses confounders through interventional pseudo-correlation augmentation.
TABLE III: Augmentation technique comparison on training U-Net. The top three results are indicated in red, magenta, and blue.
@@ IV-C Augmentation comparison
CutOut [20] 0.502 0.189 0.606 0.411 0.427
CutMix [21] 0.569 0.390 0.609 0.665 0.558
ACVC [12] 0.609 0.613 0.727 0.701 0.663
SLAug [13] 0.569 0.580 0.635 0.646 0.608
GIN-IPA [14] 0.580 0.647 0.674 0.694 0.649
FreeSDG [15] 0.632 0.671 0.677 0.690 0.668
@@
In our previous study, FreeSDG [15] utilized Gaussian filters to introduce LFS variations into samples, and combined with patch masks to merge the filtered results. However, although randomization was introduced through the augmentation of FreeSDG [15], the single-source domain, limited diversity was observed in Fig. 7 (f), resulting in suboptimal performance as indicated in Table III. Fortunately, the augmentation approach proposed in RaffeSDG successfully introduces a wide range of visual variations in the single-source domain, as illustrated in Fig. 7 (h). This diversification strategy has proven to be effective, surpassing alternative methods in training the U-Net model and achieving superior performance, as demonstrated in Table III.
TABLE IV: Comparison with SOTA segmentation algorithms. The top three results are indicated in red, magenta, and blue.
@@
CS-Net [23] / / 0.576 0.233 0.580 0.576 0.491 0.324 0.132 0.408 0.405 0.317
UTNet [24] / / 0.527 0.410 0.661 0.614 0.553 0.357 0.258 0.493 0.443 0.388
DAMAN [6] $\star$ 0.508 0.395 0.689 0.682 0.569 0.341 0.228 0.519 0.517 0.401
CSCADA [7] $\star$ 0.630 0.694 0.703 0.741 0.692 0.460 0.531 0.542 0.588 0.530
FACT [9] $\star$ 0.581 0.480 0.561 0.661 0.571 0.409 0.325 0.390 0.493 0.404
FedDG [10] $\star$ 0.663 0.659 0.752 0.697 0.693 0.496 0.492 0.602 0.535 0.531
@@
IV-D Comparison with segmentation algorithms
Comparisons with SOTA segmentation and SDG algorithms are conducted to validate the advantages of the proposed RaffeSDG. The visualized results are presented in Fig. 6, while quantitative evaluations are summarized in Table IV and V.
IV-D1 Comparison with segmentation algorithms
We employ 6 SOTA segmentation algorithms as baselines. Specifically, CS-Net [23] and UTNet [24] are utilized, which are vanilla segmentation algorithms for medical images trained solely on the training data. For domain adaptive segmentation algorithms, DAMAN [6] and CSCADA [7] are implemented, which are trained using both the training data and unlabeled test data. Additionally, FACT [9] and FedDG [10] are also included, which are segmentation algorithms based on domain generalization trained with multiple source domains.
The comparative algorithms are trained based on their data dependency, with the DRIVE dataset serving as the default training data. The segmentation results are summarized in Table IV. Due to the lack of specific design for handling domain shifts, CS-Net [23] and UTNet [24] trained on the DRIVE dataset experience a notable decline in performance when applied to other datasets. By incorporating unlabeled test data during model training, DAMAN [6] and CSCADA [7] bridge domain shifts, resulting in improved performance in the target domains. Furthermore, FACT [9] and FedDG [10] utilize multiple source domains to learn generalizable segmentation models, demonstrating decent performance on unseen target domains.
@@ IV-D1 Comparison with segmentation algorithms
TABLE V: Comparison with SDG algorithms.
@@
SAN-SAW [16] 0.582 0.560 0.584 0.531 0.564 0.410 0.389 0.412 0.361 0.393
SLAug [13] 0.626 0.662 0.717 0.683 0.672 0.456 0.495 0.559 0.519 0.507
GIN-IPA [14]* 0.649 0.654 0.712 0.717 0.683 0.481 0.486 0.553 0.559 0.520
FreeSDG [15] 0.669 0.696 0.732 0.736 0.708 0.503 0.533 0.578 0.582 0.549
@@
IV-D2 Comparison with SDG algorithms
Four SDG algorithms, including SAN-SAW [16], SLAug [13], GIN-IPA [14], and FreeSDG [15] are also compared and summarized in Table V.
SAN-SAW [16] incorporates normalization and whitening to learn domain-invariant and discriminative features, it faces challenges in effectively handling fine vessel structures. Consequently, SAN-SAW[16] exhibits false positive segmentation in Fig.6. SLAug [13] introduces a saliency-balancing fusion strategy for data augmentation, improving the out-of-domain segmentation based on the single-source domain. Meanwhile, GIN-IPA [14] employs global intensity non-linear augmentation to diversify appearances and interventional pseudo-correlation augmentation to address confounders. Though SLAug [13] and GIN-IPA [14] demonstrate impressive performance, the use of auxiliary generative networks increases the risk of model collapse.
By incorporating frequency variations and employing a jointly supervised segmentor, both FreeSDG [15] and RaffeSDG demonstrate exceptional performance compared to other SDG paradigms. They also offer the advantage of convenient implementation, eliminating the need for intricate calculations and generative network tuning. However, due to the limited diversity of frequency variations using Gaussian filters, FreeSDG [15] exhibits suboptimal performance when compared to RaffeSDG. In contrast, RaffeSDG introduces more flexible frequency variation and image blending strategies, enabling robust randomization within the single-source domain. This leads to superior performance in out-of-domain segmentation tasks.
IV-E Ablation studies
IV-E1 Data volume impact

### v3

@@ 3.2.2 Homologous sample blending
3.3 Structure consistency and saliency constraints
The proposed data augmentation promises RaffeSDG a substantial increase in the data diversity within the single-source domain. Subsequently, constraints on structure consistency and saliency are introduced to collaborate with the augmented data, further imposing model robustness and generalizability during training.
Given Gaussian filtering allows highlighting structure details [li2022annotation], it is incorporated to boost the generalizability of segmentation models by constraining on structure consistency and saliency. For an image $x$, the structure saliency map $\psi(x)$ is acquired by:
@@ 3.3 Structure consistency and saliency constraints
where $g$ is a Gaussian filter with the kernel of $(r,\sigma)$. $r$ and $\sigma$ are determined accordingQ to the image size following [li2022annotation].
Subsequently, to impose constraints on consistency and saliency, the structural saliency map serves as a self-supervised pretext task during model training. The self-supervisory loss is quantified across samples augmented from the identical vanilla images. The self-supervision loss is defined as:
@@
Specifically, the self-supervised pretext task is designed to reconstruct the structure saliency map from augmented samples. A coupling segmentation network is built to optimize the model through both self-supervision and segmentation supervision, as illustrated in Fig. 2 (c). The network consists of an encoder and two decoders. The encoder $\mathcal{E}$ processes the augmented samples $\bar{x}_{k}$ and shares its outputs with both the decoder $\mathcal{D}_{sel}$ and $\mathcal{D}_{seg}$, where $\mathcal{D}_{sel}$ predicts the structure saliency and $\mathcal{D}_{seg}$ predicts the segmentation mask.
The concrete implementation of the network is outlined in Algorithm 1. $\mathcal{E}$ is skip-connected to $\mathcal{D}_{sel}$ to facilitate the flow of contextual information, and attention mechanisms are employed to properly integrate the self-supervision into the segmentation task. The $l^{th}$h layer of the encoder and decoders are denoted as $\mathcal{E}^{l}$, $\mathcal{D}^{l}_{sel}$, and $\mathcal{D}^{l}_{seg}$, with corresponding outcomes $f^{l}_{\mathcal{E}}$, $f^{l}_{sel}$, and $f^{l}_{seg}$. The terms $CA$ and $SA$ represent typical channel and spatial attention respectively [woo2018cbam].
@@
where $\alpha$ balances $\mathcal{L}_{sel}$ and $\mathcal{L}_{seg}$.
4 Experiments
@@
4.1 Experiment settings
Table 1: Experimental data summary
@@ 4.1 Experiment settings
Vessel Fundus DRIVE, AVRDB, 40, 100, 28,
photography CHASEDB, STARE 20
@@
Joint Ultrasound 4 clinical datasets 1597, 956,
@@
4.1.1 Datasets
The experiments utilized four medical imaging modalities obtained from three different human tissues, as summarized in Table 1. The data for the vessel and retinal layer were publicly available, while the data for the joint cartilage were collected in collaboration with our partner, Southern University of Science and Technology Hospital, using discrepant ultrasound configurations of gain and grayscale. In order to generalize and assess segmentation models across these datasets, the ground-truth annotations were standardized. The datasets of DRIVE, HC-MS, and A were respectively employed as the single-source domain for each tissue, while the remaining datasets were treated as unseen target domains.
@@ 4.1.2 Implementation
4.2 Data augmentation configuration in RaffeSDG
@@
4.2.1 Filter types and blending strategies
To validate the configuration of the proposed data augmentation approach in RaffeSDG, we explored different filter types and blending strategies. The filtered and blended samples are illustrated in Fig. 4. Moreover, training data is augmented from DRIVE, and a vanilla U-Net is trained to quantitatively evaluate the performance of various configurations on the target domains, as summarized in Table 2.
@@ 4.2.1 Filter types and blending strategies
Filter type: To analyze the impact of filter types on data augmentation, we employ both the ideal filter and the Butterworth filter to introduce variations into the LFS of images. Fig.4 (b) presents the function diagram and the filtered result of the ideal filter. The sharp cutoff frequency of the ideal filter leads to the presence of ringing artifacts in the filtered result. On the other hand, the Butterworth filter offers a smoother transition between the passband and stopband, reducing the occurrence of ringing artifacts. Thus, the Butterworth filter is deemed more suitable for data augmentation, as illustrated in Fig.4 (c). Additionally, to enhance diversity in the augmented samples, we implement a channel-wise filter by applying Butterworth filters with different parameters to the RGB channels, as depicted in Fig. 4 (d). Consequently, superior performance is achieved by the channel-wise filter in Table 2.
Blending strategy: To interpret the effect of blending, we compare three strategies, including patch, grid, and continuous mask, as illustrated in Fig.4 (e). The images blended using patch and grid masks exhibit a noticeable gap, which deviates from real-world scenarios and hampers the overall context. In contrast, the continuous mask enables a seamless context in the augmented image, resulting in a more coherent representation that aligns well with real-world scenarios. The evaluation results provided in Table2 further validate our choice of filter and blending configurations. Accordingly, the channel-wise filter and the continuous mask are cooperated to implement RaffeSDG.
Table 2: Augmentation configuration comparison on training U-Net using only DRIVE for out-of-domain segmentation.
@@ 4.2.2 Parameter range
4.3 Augmentation comparison
The comparison includes both commonly used augmentation techniques and augmentation approaches designed in SDG algorithms. Specifically, the augmentation techniques include geometric transformation, CutOut [devries2017improved], and CutMix [yun2019cutmix]. And the SDG augmentation approaches are collected from ACVC [cugu2022attention], SLAug [su2022rethinking], GIN-IPA [ouyang2022causality], VFT [hu2024domain], CCDR [liang2024single], and FreeSDG [li2023frequency]. In the comparison, the dataset of DRIVE is utilized as the single-source domain, and a vanilla U-Net serves as the unified segmentor, which is trained with augmented data and conducts out-of-domain inference. The augmented samples are visualized in Fig. 6, while the segmentation performance is quantified in Table 3 by training a U-Net.
As depicted in Fig. 6, the native augmentation strategy of geometric transformation, CutOut [devries2017improved], and CutMix [yun2019cutmix] are constrained to the single-source domain, leading to limited appearance variations in the augmented samples. In contrast, the augmentation approaches used in SDG tasks present remarkable performance in Table 3. For implementing SDG in classification, ACVC [cugu2022attention] introduces 19 visual corruptions and 3 Fourier-based corruptions to augment the single-source domain. For medical image segmentation, SLAug [su2022rethinking] introduces a saliency-balancing fusion strategy that combines global-scale and local location-scale aspects to augment the single-source domain. GIN-IPA [ouyang2022causality] proposes a causality-inspired data augmentation approach, which utilizes global intensity non-linear augmentation to diversify the appearances of medical images and addresses confounders through interventional pseudo-correlation augmentation. VFT [hu2024domain] performs randomized image reconstruction with cascaded autoencoders, using a dimensionally matched intermediate output as an information-complete augmentation to enrich the source domain. Moreover, CCDR [liang2024single] combines random convolution–based texture augmentation and randomized instance normalization–based style augmentation to enrich source domain feature patterns from multiple aspects.
@@ 4.3 Augmentation comparison
Figure 6: Samples augmented by various augmentation techniques.
@@
Table 3: Augmentation technique comparison on training U-Net. The top three results are indicated in red, magenta, and blue.
@@
CutOut [devries2017improved] 0.502 0.189 0.606 0.411 0.427
CutMix [yun2019cutmix] 0.569 0.390 0.609 0.665 0.558
ACVC [cugu2022attention] 0.609 0.613 0.727 0.701 0.663
SLAug [su2022rethinking] 0.569 0.580 0.635 0.646 0.608
GIN-IPA [ouyang2022causality] 0.580 0.647 0.674 0.694 0.649
VFT [hu2024domain] 0.634 0.636 0.685 0.685 0.660
CCDR [liang2024single] 0.610 0.642 0.701 0.663 0.654
FreeSDG [li2023frequency] 0.632 0.671 0.677 0.690 0.668
@@
FreeSDG [li2023frequency] utilized Gaussian filters and patch masks to augment the single-source domain. However, in Fig. 6 (f), limited diversity was observed in the augmentation performed by FreeSDG [li2023frequency], resulting in suboptimal performance as indicated in Table 3. Fortunately, the augmentation approach proposed in RaffeSDG successfully introduces a wide range of visual variations in the single-source domain, as illustrated in Fig. 6 (h). This diversification strategy has proven to be effective, surpassing alternative methods in training the U-Net model and achieving superior performance, as demonstrated in Table 3.
Table 4: Comparison with SOTA segmentation algorithms. The top three results are indicated in red, magenta, and blue.
@@
CS-Net [mou2019cs] / / 0.576 0.233 0.580 0.576 0.491 0.324 0.132 0.408 0.405 0.317
UTNet [gao2021utnet] / / 0.527 0.410 0.661 0.614 0.553 0.357 0.258 0.493 0.443 0.388
Dtmformer [wang2024dtmformer] / / 0.618 0.628 0.714 0.683 0.661 0.484 0.458 0.573 0.520 0.501
Rolling-unet [liu2024rolling] / / 0.622 0.598 0.713 0.651 0.646 0.456 0.421 0.567 0.483 0.482
FDA [yang2020fda] $\star$ 0.660 0.683 0.759 0.720 0.705 0.495 0.514 0.618 0.564 0.548
DAMAN [mukherjee2022domain] $\star$ 0.508 0.395 0.689 0.682 0.569 0.341 0.228 0.519 0.517 0.401
CSCADA [gu2022contrastive] $\star$ 0.630 0.694 0.703 0.741 0.692 0.460 0.531 0.542 0.588 0.530
AIF-SFDA [li2025aif] $\star$ 0.670 0.648 0.762 0.719 0.700 0.508 0.481 0.622 0.562 0.543
FACT [xu2021fourier] $\star$ 0.581 0.480 0.561 0.661 0.571 0.409 0.325 0.390 0.493 0.404
FedDG [liu2021feddg] $\star$ 0.663 0.659 0.752 0.697 0.693 0.496 0.492 0.602 0.535 0.531
DDG-Med [cheng2025dynamic] $\star$ 0.661 0.668 0.751 0.672 0.688 0.495 0.503 0.608 0.507 0.528
@@
4.4 Comparison with segmentation algorithms
Comparisons with SOTA segmentation and SDG algorithms are conducted to validate the advantages of the proposed RaffeSDG. The visualized results are presented in Fig. 7, while quantitative evaluations are summarized in Table 4 and 5.
4.4.1 Comparison with segmentation algorithms
We employ 11 SOTA segmentation algorithms as baselines. Specifically, CS-Net [mou2019cs], UTNet [gao2021utnet], Dtmformer [wang2024dtmformer] and Rolling-unet [liu2024rolling] are utilized, which are vanilla segmentation algorithms for medical images trained solely on the training data. For domain adaptive segmentation algorithms, FDA [yang2020fda], DAMAN [mukherjee2022domain], CSCADA [gu2022contrastive] and AIF-SFDA [li2025aif] are implemented, which are trained using both the training data and unlabeled test data. Additionally, FACT [xu2021fourier], FedDG [liu2021feddg] and DDG-Med [cheng2025dynamic] are also included, which are segmentation algorithms based on domain generalization trained with multiple source domains.
The comparative algorithms are trained based on their data dependency, with the DRIVE dataset serving as the default training data. The segmentation results are summarized in Table 4. Due to the lack of specific design for handling domain shifts, CS-Net [mou2019cs], UTNet [gao2021utnet], Dtmformer [wang2024dtmformer] and Rolling-unet [liu2024rolling] trained on the DRIVE dataset experience a notable decline in performance when applied to other datasets. By incorporating unlabeled test data during model training, FDA [yang2020fda], DAMAN [mukherjee2022domain], CSCADA [gu2022contrastive] and AIF-SFDA [li2025aif] bridge domain shifts, resulting in improved performance in the target domains. Furthermore, FACT [xu2021fourier], FedDG [liu2021feddg] and DDG-Med [cheng2025dynamic] utilize multiple source domains to learn generalizable segmentation models, demonstrating decent performance on unseen target domains.
@@ 4.4.1 Comparison with segmentation algorithms
Table 5: Comparison with SDG algorithms.
@@
SLAug [su2022rethinking] 0.626 0.662 0.717 0.683 0.672 0.456 0.495 0.559 0.519 0.507
GIN-IPA [ouyang2022causality]* 0.649 0.654 0.712 0.717 0.683 0.481 0.486 0.553 0.559 0.520
VFT [hu2024domain] 0.638 0.646 0.705 0.710 0.675 0.470 0.473 0.546 0.551 0.510
CCDR [liang2024single] 0.652 0.706 0.780 0.705 0.711 0.484 0.547 0.643 0.544 0.554
LRFS [arslan2024single] 0.625 0.621 0.638 0.646 0.632 0.464 0.459 0.476 0.483 0.471
StyDeSty [liu2024stydesty] 0.618 0.627 0.637 0.641 0.631 0.448 0.456 0.466 0.471 0.460
FreeSDG [li2023frequency] 0.669 0.696 0.732 0.736 0.708 0.503 0.533 0.578 0.582 0.549
@@
4.4.2 Comparison with SDG algorithms
Seven SDG algorithms, including SLAug [su2022rethinking], GIN-IPA [ouyang2022causality], VFT [hu2024domain], CCDR [liang2024single], LRFS [arslan2024single], StyDeSty [liu2024stydesty], and FreeSDG [li2023frequency] are also compared and summarized in Table 5. SLAug [su2022rethinking] introduces a saliency-balancing fusion strategy for data augmentation, which enhances out-of-domain segmentation performance under a single-source domain setting. GIN-IPA [ouyang2022causality] employs global intensity non-linear augmentation to diversify image appearances and utilizes interventional pseudo-correlation augmentation to mitigate the influence of confounding factors. VFT [hu2024domain] enhances model generalization by integrating Hessian matrix-based vector flow feature extraction with latent space uncertainty-guided image augmentation. CCDR [liang2024single] adopts domain randomization in terms of texture and style and performs class-based domain alignment between the source and auxiliary domains to improve adaptation to the target domain. LRFS [arslan2024single] enforces a Lipschitz constraint in the image frequency spectrum, reducing the sensitivity to high-frequency components and improving the robustness in the mid-frequency range, thus lowering the reliance on domain-specific information. Meanwhile, StyDeSty [liu2024stydesty] utilizes a min-max cooperative optimization framework between stylization and de-stylization modules, enabling the model to learn domain-invariant representations. Though GIN-IPA [ouyang2022causality] and CCDR [liang2024single] demonstrate impressive performance, the use of auxiliary generative networks increases the risk of model collapse.
By incorporating frequency variations and employing a jointly supervised segmentor, both FreeSDG [li2023frequency] and RaffeSDG demonstrate exceptional performance compared to other SDG paradigms. They also offer the advantage of convenient implementation, eliminating the need for intricate calculations and generative network tuning. However, due to the limited diversity of frequency variations using Gaussian filters, FreeSDG [li2023frequency] exhibits suboptimal performance when compared to RaffeSDG. In contrast, RaffeSDG introduces more flexible frequency variation and image blending strategies, enabling robust randomization within the single-source domain. This leads to superior performance in out-of-domain segmentation tasks.
4.5 Ablation studies
4.5.1 Data volume impact